Please help me solve this equation (5x^2+7x+8)-(2x^2+3x+4)
Answer by checkley75(3666)

5x^2+7x+8
-2x^2-3x-4
----------------------
3x^2+4x+4
(3x-2)(x+2)=0
3x-2=0
3x=2
x=2/3 answer
x+2=0
x=-2 answer